OpenStudy (anonymous):

Find X 10^(2x) + 10^x = 20

OpenStudy (freckles):

let 10^x=u you have a quadratic equation to solve

OpenStudy (freckles):

\[u^2+u=20\]

OpenStudy (freckles):

See if you can solve that for u

OpenStudy (anonymous):

\[u^{2} = 20- u\]

OpenStudy (freckles):

\[u^2+u-20=0\]

OpenStudy (freckles):

Next you can factor u^2+u-20 or you can use the quadratic formula next

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

wait but that still doesn't help me solve for x

OpenStudy (freckles):

yes it does what did you get for u

OpenStudy (freckles):

remember u is 10^x

OpenStudy (anonymous):

4 and -5

OpenStudy (freckles):

when we are done solving for u we can replace u with 10^x and then solve for x

OpenStudy (anonymous):

ok

OpenStudy (freckles):

I think you have it backwards

OpenStudy (freckles):

\[(10^x)^2+10^x-20= \\ (10^x+5)(10^x-4)=0 \\ \] lol no you are right I looked at the wrong end of my paper

OpenStudy (freckles):

\[10^x=-5 \text{ or } 10^x=4 \]

OpenStudy (freckles):

Now can the first equation ever happen over the real numbers x?

OpenStudy (anonymous):

where did the square go?

OpenStudy (freckles):

what square?

OpenStudy (anonymous):

the (10^x)^2

OpenStudy (freckles):

\[u^2+u-20 \\ (u+5)(u-4) \\ \text{ remember } u \text{ is } 10^x \]

OpenStudy (anonymous):

oh, thats right

OpenStudy (freckles):

10^x *10^x=10^(x+x)=10^(2x)

OpenStudy (anonymous):

so what do we do now?

OpenStudy (anonymous):

(u+5)(u-4) = 0

OpenStudy (freckles):

well I thought you already solved for u... you are going backwards now

OpenStudy (freckles):

you had u=-5 or u=4 remember to replace u with 10^x

OpenStudy (freckles):

I gave those equations above some place

OpenStudy (freckles):

\[10^x=-5 \text{ or } 10^x=4 \]

OpenStudy (freckles):

now I was asking can that first equation ever happen?